Ttiangle ABC has sides of length 6cm, 8cm, and 12cm. Triangle EFG is similar to triangle ABC, which means the sides of the two triangles are proportional. The shortest side of triangle EFG is 9cm in length. What is the perimeter of triangle EFG?
Answer by stanbon(48558)

If shortest of one triangle to shortest of the other is 6 to 9 or 1 to 1.5
the same ratio holds for the missing sides of the larger triangle.
So the larger trianglehas sides of (1.5)*8=12 and (1/5)*12=18
so the perimeter of the larger triangle is 9+12+18=39
--------------
Cheers,
Stan H.